OpenStudy (anonymous):

so i just need someone too check something for me In the equation y equals one-fourth x plus 8, what is y when x is twenty? 13 is what i got

OpenStudy (anonymous):

@KingGeorge can u check this for me

OpenStudy (mertsj):

\[y=\frac{1}{4}x+8\] \[y=\frac{1}{4}(\frac{20}{1})+8\] \[y=5+8=13\]

OpenStudy (nurali):

y=1/4 x 20 + 8 Remove the parentheses from the numerator. y=(20)/(4)+8 Reduce the expression (20)/(4) by removing a factor of 4 from the numerator and denominator. y=5+8 Add 8 to 5 to get 13. y=13

OpenStudy (anonymous):

well i have this and i got 2 Here is a table of values for the linear equation y equals 2 x minus 4. The x value is 3. The y value is missing. What is the missing number? nineteen twenty-four ten 2

OpenStudy (nurali):

2x-4. 2*3-4 6-4 2

OpenStudy (anonymous):

thank yall

OpenStudy (mertsj):

yw

OpenStudy (nurali):

Anytime.